OpenStudy (anonymous):

Let h=v0^2/4.9 sin(theta)cos(theta) model the horizontal distance in meters traveled by a projectile. If the initial velocity is 36 meters/second, which equation would you use to find the angle needed to travel 100 meters? 7.36sin(2) = 100 132.12sin(2) = 100 264.49sin(2) = 100 100sin(2) = 100

OpenStudy (tylerd):

the the v0 the initial velocity?

OpenStudy (anonymous):

behold the most helpful tool on earth

OpenStudy (anonymous):

The initial velocity is 36

OpenStudy (tylerd):

\[h=\frac{ v _{0}^2 }{ 4.9\sin(\theta)\cos(\theta) }\]

OpenStudy (tylerd):

does it look like that?

OpenStudy (anonymous):

no the sin and cos are to the right. they are being multiplied, not divided

OpenStudy (tylerd):

\[h=v _{0}^24.9\sin(\theta)\cos(\theta)\]

OpenStudy (tylerd):

?

OpenStudy (anonymous):

sorry. idk how to send pictures, but everything was right before just the sin and cos were being multiplied to the (v0^2)/4.9

OpenStudy (anonymous):

so basicall ((vo^2)/4.9)sin(theta)cos(theta)

OpenStudy (tylerd):

ok so v0=36m/s

OpenStudy (anonymous):

the equation isnt right...

OpenStudy (tylerd):

browser keeps acting up...

OpenStudy (tylerd):

\[\frac{ v_{0}^2\sin(\theta)\cos(\theta) }{ 4.9 }\]

OpenStudy (anonymous):

did you get it?

OpenStudy (tylerd):

sin(2x)=2sin(x)cos(x) so sin(x)cos(x)=(1/2)sin(2x)

OpenStudy (anonymous):

ok

OpenStudy (tylerd):

ok looks like 36^2/4.9 = 264.49

OpenStudy (tylerd):

so we have \[(264.49)(\frac{ 1 }{ 2 })(\sin(2x))=100\]

OpenStudy (anonymous):

ohh ok

OpenStudy (tylerd):

cause i think sin(2x)=2sin(x)cos(x) and we have sin(x)cos(x) so we divide the 2 out

OpenStudy (tylerd):

id say im 80% sure on this one but its been awhile

OpenStudy (anonymous):

i see. ok. thanks for the help

OpenStudy (anonymous):

wait so if do i multiply the 264.49 by1/2?

OpenStudy (tylerd):

ya i got something like 132.245

OpenStudy (tylerd):

and the only answer chose is 132.12

OpenStudy (tylerd):

so not sure on that

OpenStudy (anonymous):

oh ok. thanks

OpenStudy (tylerd):

yep